OpenStudy (anonymous):

slope m= ? -6x+3y=0 can some help x,y for my graph. tks

OpenStudy (anonymous):

rearrange your equation into general form y=mx+c which m is the slope.

OpenStudy (anonymous):

I dont understand, cause it written then we have to follow the formula but it doesnt come out 6\[-6_{0}+3y=0\] x y 0 -6 I'm just confused as to what formula to use so I can have coordinates to point on the graph

OpenStudy (anonymous):

As melor said, write the equation in the form y=mx+c. m then will be the slope. To do so, first move 6x to the other side, you get: \[3y=6x\] Now, divide both sides by 3, you will have:\[{3y \over 3}={6x \over 3} \implies y=2x\] So the equation is y=2x+0. What is the slope?

OpenStudy (anonymous):

its -6x+3y=0

OpenStudy (anonymous):

I know. I just rearranged the equation to the form y=2x. The slope is now the coefficient of x, which is 2.

OpenStudy (anonymous):

-6x+3y=0, bring -6x to the right hand side, 3y=6x, divide all by 3, y=2x. compare what you have with the general form y=mx+c, m=2, y=0 since there is no terms after x :)

OpenStudy (anonymous):

sorry, not y=0, it's c=0 :) which c is the y-intercept.

OpenStudy (anonymous):

tks i'm going try using the formula

OpenStudy (anonymous):

you should be familiar with the general form of a straight line, it helps you alot in searching for gradient or y-intercept :)